OpenStudy (anonymous):

the quotient of N and -0.5 is 13 WHAT IS THE VALUE ON N CAN U SHOW WORK PLEASE

OpenStudy (anonymous):

The "quotient of N and -0.5 is 13" means that N divided by -0.5 is equal to 13. \[N \div (-0.5) = 13\] Solve for N by multiplying both sides of the equation by -0.5. \[ N \div (-0.5) \times (-0.5) = 13 \times (-0.5) \] \[N = \frac{-13}{2} = -6.5 \]

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

You're welcome